# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ 1 Introduction
It is widely believed that new physics is required to explain various phenomena beyond the Standard Model (SM), including the neutrino mass and mixing King:2014nza , dark matter Bertone:2016nfn , strong CP problem Marsh:2015xka , and matter-antimatter asymmetry Dine:2003ax , etc. As the cosmic temperature drops with the expansion of the Universe, the spontaneous breakdown of discrete symmetries during a phase transition in new physics models can generate sheet-like topological defects, the domain walls Zeldovich:1974uw ; Kibble:1976sj . The Universe after the phase transition is then randomly divided by them into distinct domains, which represent the clumps of the degenerate vacua. The domain walls tend to stretch in size and the neighboring domains with the same vacua merge to stabilize the network. This evolution can be described by the percolation theory Stauffer:1978kr , whose fundamental results are based on the existence of a percolation threshold $p_{c}=0.311$ for a simple cubic lattice. The domain wall’s energy density scales as $\rho_{w}\propto t^{-1}$ in the scaling regime, which is much slower than the decay of radiation and dust, and eventually dominates the cosmological evolution. In the domain wall era, the repulsive force from domain walls would lead to an inflation phase of the Universe with the scale factor going as $a\propto t^{2}$. Such an expansion history is not compatible with the observations since it could strongly alter the Big Bang nucleosynthesis (BBN), reduce the time for the galaxy formation, and lead to unacceptably large anisotropies of the cosmic microwave background (CMB) Planck:2015fie . Nevertheless, the research on the nature of dark energy in the context of domain wall has been continuously conducted by recent works Friedland:2002qs ; Conversi:2004pi ; Mulki:2022tfw .
@@
Discrete symmetries exist widely in new physics models. A $Z_{2}$ symmetry can emerge from the spontaneous breakdown of a gauge group in the grand unified theory (GUT) Preskill:1992ck ; Dunsky:2021tih . The $Z_{2}$ symmetry is also commonly imposed on the two-Higgs-doublet models for the spontaneous CP violation to suppress the Higgs-mediated flavor-changing neutral currents (FCNCs) at tree level Lee:1973iz ; Preskill:1991kd ; Gunion:2002zf . The global $U(1)$ symmetry in the axion models for solving the strong CP problem is explicitly broken to the $Z_{\mathcal{N}}$ symmetry by the QCD instanton effect Sikivie:1982qv ; Preskill:1991kd . Furthermore, to provide a dark matter candidate, a $Z_{2}$ symmetry is often imposed to make the dark matter particle stable Chao:2017vrq ; Huang:2017kzu ; Grzadkowski:2018nbc ; Chiang:2019oms ; Chiang:2020yym ; King:2023ayw ; Pham:2024vso .
@@
A common solution to the domain wall problem is to introduce a bias potential $V_{\rm bias}$ to break the degenerate vacua Dine:1993yw ; Dvali:1994wv . The domain wall network collapses as its surface energy is comparable with the bias potential, $\sigma H\simeq V_{\rm bias}$. The bias potential can be generated by the instanton effects if the global symmetry is anomalous under the framework of the theory, with such examples as the QCD and quantum gravity Preskill:1991kd ; Dine:1993yw ; Dvali:1994wv ; Abel:1995wk ; Larsson:1996sp ; Lu:2023mcz . It is notable that based on the successful and precise prediction in Ref. Chiang:2020aui , Ref. Lu:2023mcz suggest that the recent pulsar timing array (PTA) observations, including the NANOGrav 15-year dataset (NG15) and IPTA-DR2 dataset, strongly support the mechanism of the QCD instanton-induced domain wall annihilations.
@@
Primordial black holes (PBHs) have attracted a lot of attention since it was first suggested by Zel’dovich more than 50 years ago Zeldovich:1967lct . Contrary to the black holes resulting from the death and collapse of stars in their final evolution stage, PBHs are formed by cosmological processes in the early Universe. In 1971, Hawking proposed the first mechanism of PBH production Hawking:1971ei ; Carr:1974nx : the fluctuations of a radiation-dominated Universe can produce some overdense regions. Once the density of the region $\delta=\Delta\rho/\bar{\rho}$ exceeds a critical value $\delta_{c}\simeq w$, the overdense region can stop expanding and re-collapse to form PBHs when the region enters the horizon. Cosmic inflation Ivanov:1994pa ; Randall:1995dj ; Garcia-Bellido:1996mdl is commonly considered a primary source of density perturbation. The large curvature perturbations generated by the quantum fluctuations of the scalar fields are dominated by the density perturbations once they exit the horizon Garcia-Bellido:1996mdl . Other sources for PBH production include the collapse of topological defects such as cosmic strings Hawking:1987bn ; Polnarev:1988dh ; Brandenberger:2021zvn and closed domain walls Ipser:1983db ; Ferrer:2018uiu ; Liu:2019lul ; Gouttenoire:2023gbn ; Ge:2023rrq , the density perturbation from Q-balls during a matter-dominated era Cotner:2016cvr ; Cotner:2019ykd , the collision of bubble walls Crawford:1982yz ; Kodama:1982sf ; Hawking:1982ga ; Johnson:2011wt ; Kusenko:2020pcg and the collapse of false-vacuum patches during a first-order phase transition Baker:2021sno ; Gross:2021qgx ; Kawana:2021tde ; Liu:2021svg ; Baldes:2023rqv ; Flores:2024lng ; Gouttenoire:2023naa ; Lewicki:2023ioy .
@@
In contrast to PBH formation from the collapse of closed domain wall Ipser:1983db ; Ferrer:2018uiu ; Liu:2019lul ; Gouttenoire:2023gbn ; Ge:2023rrq , we show in this work that the density perturbation due to the number density fluctuations of horizon-size domain walls in the late stage could lead to the formation of PBHs. The collapse of a closed domain wall needs to satisfy the Schwarzschild criterion, i.e., the closed wall has a Schwarzschild radius near collapse Ferrer:2018uiu . Our scenario is based on the critical collapse of horizon-size perturbations, as outlined above. We find that if the annihilations of the domain wall are the source of the nano-Hz stochastic gravitational wave background (SGWB) observed recently by PTAs, a significant PBH relic abundance could be produced via the number density fluctuations of the walls before their annihilations.
@@
This work is arranged as follows. In section 2, we briefly review the equation of state and evolution of the domain wall. In section 3, we calculate the variance of density contrast from domain wall number density fluctuations. In section 4, we review the critical collapse mechanism. In section 5, we calculate the PBH relic abundance from domain wall number density fluctuations. In section 6, we define an effective equation of state for the mixture gas containing domain walls and radiation. A fit formula for the $\delta_{c}-\omega_{\rm eff}$ relation from the simulation is also provided. We derive the gravitational wave background from the domain wall annihilations in section 7, and perform the Bayesian statistics analysis of the IPTA-DR2 and NG15 datasets in section 8. In section 9, We show the constraints on the domain wall interpretations for the PTAs using the PBH abundance from the domain wall number density fluctuations. Finally, in section 10, we summarize our conclusions and discuss the uncertainty in the PBH formations. Various particle physics models for the realization of $N$ domain walls for PTAs are also discussed in this section. In appendix A, appendix B, and appendix C, we provide respectively the details of calculations for the energy-momentum tensor of the domain wall in a $Z_{2}$ symmetry model, the surface energy tensor in the observer frame, and the velocity-dependent one-scale (VOS) model. In appendix D, we provide details for the complementary error function. In appendix E, we present further details of our numerical strategy for the PTA dataset.
@@ 2.1 Domain wall equation of state
The energy-momentum tensor of a plane domain wall lying in the $yz$ plane at $x=0$ (neglecting the thickness of the wall) can be estimated as Vilenkin:1984ib ; Gelmini:1988sf
$T_{\mu\nu}(x)=\sigma_{w}\delta(x){\rm diag}(1,0,-1,-1),$ (1)
where $\sigma_{w}$ is the surface tension of the wall. In appendix A, we provide a concrete example for the calculation of the energy-momentum tensor in a toy model with a $Z_{2}$ symmetry. Consider that within a volume $V\equiv L^{3}$, there are $N$ $yz$-plane walls moving along the $x$ direction with an average velocity $\overline{v}_{w}$. The average energy-momentum tensor within a length $L$ can be estimated as
[…]
@@
[…]
with the superscript $x$ denoting the axis of wall motion. The energy-momentum tensor (1) is in the rest frame of the wall. We can obtain the energy-momentum tensor in the observer frame by performing a Lorentz transformation. In general, the domain wall can move along an arbitrary direction in the three-dimensional space and, with some algebra, we obtain the surface energy-momentum tensor in the observer frame as
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2),$ (4)
@@
$\rho_{w}=\overline{T}_{00}=\frac{\gamma^{2}\sigma_{w}}{\overline{L}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ p_{w}=\overline{T}_{ii}=\frac{\sigma_{w}}{3\overline{L}}(\overline{v}_{w}^{2}\gamma^{2}-2).$ (5)
@@
$\omega=\frac{p_{w}}{\rho_{w}}=\frac{-2+\overline{v}_{w}^{2}\gamma^{2}}{3\gamma^{2}}.$ (6)
@@ 2.2 The evolution of a domain wall
The correlation length $\overline{L}$ and velocity $\overline{v}_{w}$ are two main physics quantities for describing the domain walls. The evolution of $\overline{L}$ and $\overline{v}_{w}$ with the cosmic expansion can be described by the VOS model Martins:2016book (see appendix C for more details)
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (7)
@@
where the momentum parameter $k_{w}=0.66\pm 0.04$ and the chopping parameter $c_{w}=0.81\pm 0.04$ from the simulations of domain wall evolution in a radiation-dominated universe Martins:2016ois . The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}}.$ (8)
@@
The first term on the right-hand side of Eq. (8) describes the Hubble damping effect due to the cosmic expansion and the second term denotes the particle friction, which is related to the particle pressure by $\ell_{f}=\overline{v}_{w}\sigma_{w}/\Delta P$, where the pressure between the wall and particles $\Delta P$ is model-dependent.
Figure 1: The evolution of $\overline{L}$ (red curves) and $\overline{v}_{w}$ (blue curves) as a function of cosmic time $t/t_{s}$. The results without any particle friction are denoted by the solid lines. From the upper left to the lower right, the dot-dashed lines represent the results with $N_{c}=1$, 5, 15, and 30, respectively.
In Fig. 1, we depict the evolution of the domain wall formed at time $t_{s}$ with a symmetry-breaking scale $v=\sigma_{w}^{1/3}=10^{5}$ GeV. We also assume a pressure induced by friction from the interaction between wall and QCD gluons given by
$\Delta P=\frac{24N_{c}^{2}\alpha_{s}^{2}}{\pi^{4}}\overline{v}_{w}T^{4},$ (9)
where we will adopt the strong coupling $\alpha_{s}=1$ for illustration and note that $\alpha_{s}\lesssim 0.1$ for the energy scale at $\sim 10^{5}$ GeV, and the strength of the pressure is controlled by the parameter $N_{c}$. For the solid curves, the friction is removed by setting $N_{c}=0$. From the upper left panel to the lower right panel of Fig. 1, the dot-dashed curves represent the results with $N_{c}=1$, 5, 15, and 30, respectively. The results without particle friction indicate that after its formation, the domain wall rapidly expands to a superhorizon size, $\overline{L}\sim 6t$, due to the tension of the wall. At time $t\sim 20t_{s}$, the Hubble friction from the cosmic expansion becomes significant and damps the further expansion of the wall. The evolution of the domain wall finally enters into the scaling regime, in which the domain wall is of the horizon size, $\overline{L}\simeq t$, and the velocity of the wall $\overline{v}_{w}\simeq 0.4$. For common models with $N_{c}=1$, the friction effect from interaction with gluons is negligible. When $N_{c}\gtrsim 5$ the suppression from particle friction on the domain wall growing at $t\lesssim 20t_{s}$ becomes significant. The suppression effect stops improving much for $N_{c}\gtrsim 15$. A nontrivial conclusion from Fig. 1 is that the sizable friction (if existing in the model) from the surrounding plasma can not stop the walls from eventually entering into the scaling regime (for $t\gtrsim 10^{4}t_{s}$). Therefore, the mechanism of the PBH production from domain wall number density fluctuations is independent of the particle physics models.

### v2

# Probing Primordial Black Hole Formation from Domain Wall Isocurvature Perturbations: Constraints and Implications

@@ 1 Introduction
It is widely believed that new physics is required to explain various phenomena beyond the Standard Model (SM), including the neutrino mass and mixing [1], dark matter [2], strong CP problem [3], and matter-antimatter asymmetry [4], etc. As the cosmic temperature drops with the expansion of the Universe, the spontaneous breakdown of discrete symmetries during a phase transition in new physics models can generate sheet-like topological defects, the domain walls [5, 6]. The Universe after the phase transition is then randomly divided by them into distinct domains, which represent the clumps of the degenerate vacua. The domain walls tend to stretch in size and the neighboring domains with the same vacua merge to stabilize the network. This evolution can be described by the percolation theory [7], whose fundamental results are based on the existence of a percolation threshold $p_{c}=0.311$ for a simple cubic lattice. The domain wall’s energy density scales as $\rho_{w}\propto t^{-1}$ in the scaling regime, which is much slower than the decay of radiation and dust, and eventually dominates the cosmological evolution. In the domain wall era, the repulsive force from domain walls would lead to an inflation phase of the Universe with the scale factor going as $a\propto t^{2}$. Such an expansion history is not compatible with the observations since it could strongly alter the Big Bang nucleosynthesis (BBN), reduce the time for the galaxy formation, and lead to unacceptably large anisotropies of the cosmic microwave background (CMB) [8]. Nevertheless, the research on the nature of dark energy in the context of domain wall has been continuously conducted by recent works [9, 10, 11].
@@
Discrete symmetries exist widely in new physics models. A $Z_{2}$ symmetry can emerge from the spontaneous breakdown of a gauge group in the grand unified theory (GUT) [12, 13]. The $Z_{2}$ symmetry is also commonly imposed on the two-Higgs-doublet models for the spontaneous CP violation to suppress the Higgs-mediated flavor-changing neutral currents (FCNCs) at tree level [14, 15, 16]. The global $U(1)$ symmetry in the axion models for solving the strong CP problem is explicitly broken to the $Z_{\mathcal{N}}$ symmetry by the QCD instanton effect [17, 15]. Furthermore, to provide a dark matter candidate, a $Z_{2}$ symmetry is often imposed to make the dark matter particle stable [18, 19, 20, 21, 22, 23, 24].
@@
A common solution to the domain wall problem is to introduce a bias potential $V_{\rm bias}$ to break the degenerate vacua [25, 26]. The domain wall network collapses as its surface energy is comparable with the bias potential, $\sigma H\simeq V_{\rm bias}$. The bias potential can be generated by the instanton effects if the global symmetry is anomalous under the framework of the theory, with such examples as the QCD and quantum gravity [15, 25, 26, 27, 28, 29]. It is notable that based on the successful and precise prediction in Ref. [30], Ref. [29] suggest that the recent pulsar timing array (PTA) observations, including the NANOGrav 15-year dataset (NG15) and IPTA-DR2 dataset, strongly support the mechanism of the QCD instanton-induced domain wall annihilations.
@@
Primordial black holes (PBHs) have attracted much attention since it was first suggested by Zel’dovich more than 50 years ago [31]. Contrary to the black holes resulting from the death and collapse of stars in their final evolution stage, PBHs are formed by cosmological processes in the early Universe. In 1971, Hawking proposed the first mechanism of PBH production [32, 33]: the fluctuations of a radiation-dominated Universe can produce some overdense regions. Once the density of the region $\delta=\delta\rho/\bar{\rho}$ exceeds a critical value $\delta_{c}\simeq w$, the overdense region can stop expanding and re-collapse to form PBHs when the region enters the horizon. Cosmic inflation [34, 35, 36] is commonly considered a primary source of density perturbation. The large curvature perturbations generated by the quantum fluctuations of the scalar fields are dominated by the density perturbations once they exit the horizon [36]. For other PBH production mechanisms, we refer to Refs. [37, 38] for a recent review.
@@
In contrast to PBH formation from the collapse of closed domain wall [39, 40, 41, 42, 43, 44], we show in this work that the density perturbation due to the number density fluctuations of horizon-size domain walls in the late stage could lead to the formation of PBHs. The collapse of a closed domain wall needs to satisfy the Schwarzschild criterion, i.e., the closed wall has a Schwarzschild radius near collapse [41]. Our scenario is based on the critical collapse of horizon-size perturbations, as outlined above. We find that if the annihilations of the domain wall are the source of the nano-Hz stochastic gravitational wave background (SGWB) observed recently by PTAs, a significant PBH relic abundance could be produced via the number density fluctuations of the walls before their annihilations.
@@
This work is arranged as follows. In section 2, we review the evolution of the domain wall and emphasize the chopping effect on the domain wall size. In section 3, we calculate the variance of density contrast and the power spectrum of the density perturbation from domain wall number density fluctuations. In section 5, we review the critical collapse mechanism. In section 6, we calculate the PBH relic abundance from domain wall number density fluctuations. In section 7, we define an effective equation of state for the mixture gas containing domain walls and radiation. A fit formula for the $\delta_{c}-\omega_{\rm eff}$ relation from the simulation is also provided. We derive the gravitational wave background from the domain wall annihilations in section 8, and perform the Bayesian statistics analysis of the IPTA-DR2 and NG15 datasets in section 9. In section 10, We show the constraints on the domain wall interpretations for the PTAs using the PBH abundance from the domain wall number density fluctuations. Finally, in section 11, we summarize our conclusions and discuss the uncertainty in the PBH formations. In appendix A, appendix B, and appendix C, we provide respectively the details of calculations for the energy-momentum tensor of the domain wall in a $Z_{2}$ symmetry model, the surface energy tensor in the observer frame, and the velocity-dependent one-scale (VOS) model. In appendix D, we provide details for the complementary error function. In appendix E, we present further details of our numerical strategy for the PTA dataset. In appendix F, we provide detailed models that can generate $N$ domain walls after spontaneous symmetry breaking.
@@ 2.1 Domain wall equation of state
The energy-momentum tensor of a plane domain wall lying in the $yz$ plane at $x=0$ (neglecting the thickness of the wall) can be estimated as [45, 46]
$T_{\mu\nu}(x)=\sigma_{w}\delta(x){\rm diag}(1,0,-1,-1),$ (2.1)
where $\sigma_{w}$ is the surface tension of the wall. In appendix A, we provide a concrete example for the calculation of the energy-momentum tensor in a toy model with a $Z_{2}$ symmetry. Consider that within a volume $V\equiv L^{3}$, there are NN yzyz-plane walls moving along the $x$ direction with an average velocity $\overline{v}_{w}$. The average energy-momentum tensor within a length $L$ can be estimated as
$\begin{aligned} \overline{T}_{\mu\nu}\simeq\frac{1}{L}\int_{-L/2}^{L/2}dxT_{\mu\nu}(x,\overline{v}_{w})=\frac{NS_{\mu\nu}^{x}(\overline{v}_{w})}{L}=\frac{S_{\mu\nu}^{x}(\overline{v}_{w})}{\overline{L}},\end{aligned}$ (2.2)
@@
$S_{\mu\nu}^{x}(\overline{v}_{w})=\int_{\overline{L}}dxT_{\mu\nu}(x,\overline{v}_{w}),$ (2.3)
with the superscript $x$ denoting the axis of wall motion. The energy-momentum tensor (2.1) is in the rest frame of the wall. We can obtain the energy-momentum tensor in the observer frame by performing a Lorentz transformation. In general, the domain wall can move along an arbitrary direction in the three-dimensional space and, with some algebra, we obtain the surface energy-momentum tensor in the observer frame as
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(3\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2),$ (2.4)
@@
$\rho_{w}=\overline{T}_{00}=\frac{\gamma^{2}\sigma_{w}}{\overline{L}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ p_{w}=\overline{T}_{ii}=\frac{\sigma_{w}}{3\overline{L}}(\overline{v}_{w}^{2}\gamma^{2}-2).$ (2.5)
@@
$\omega=\frac{p_{w}}{\rho_{w}}=\frac{-2+\overline{v}_{w}^{2}\gamma^{2}}{3\gamma^{2}}.$ (2.6)
@@ 2.2 The evolution of a domain wall
The correlation length $\bar{L}$ (or curvature radius) and the root-mean-square velocity $\bar{v}_{w}$ are two key physical quantities for describing the domain walls. We first adopt the so-called VOS model [47] (see appendix C for more details) to describe the evolution of $\bar{L}$ and $\bar{v}_{w}$ with cosmic expansion, which is given by
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (2.7)
@@
where the momentum parameter $k_{w}=0.66\pm 0.04$ and the chopping parameter $c_{w}=0.81\pm 0.04$ are derived from simulations of domain wall evolution in a radiation-dominated Universe [48]. The momentum parameter $k_{w}$ characterizes the curvature acceleration of the wall, while the chopping parameter $c_{w}$ represents the interaction and merger between walls. The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}}.$ (2.8)
@@
The first term on the right-hand side of Eq. (2.8) describes the Hubble damping effect due to the cosmic expansion and the second term denotes the particle friction, which is related to the pressure of the interaction by $\ell_{f}=\overline{v}_{w}\sigma_{w}/\Delta P$, where the pressure between the wall and particles $\Delta P$ is model-dependent. We assume a pressure induced by friction from the bias potential, expressed as $\Delta P=V_{\rm bias}\equiv V_{\rm false}-V_{\rm true}$. The vacuum pressure arising from the bias potential can drive the annihilation of the domain wall once it becomes comparable to the pressure exerted by the wall’s surface tension. This condition determines the annihilation time of the domain wall as follows
$t_{\rm ann}=\frac{C_{d}\mathcal{A}\sigma_{w}}{V_{\rm bias}},$ (2.9)
where $\mathcal{A}$ is an $\mathcal{O}(1)$ parameter that depends on the number of domains.
Figure 1: Upper panel: The evolution of $\bar{L}$ (red curves) and $\bar{v}_{w}$ (blue curves) as a function of cosmic time $t/t_{s}$ in the VOS model is shown. The solid curves and the dot-dashed curves represent the results with a chopping parameter $c_{w}=0.81$ and $0.61$, respectively. Lower panel: The evolution of the closed domain wall, governed by Eq. (2.10), is illustrated for different initial radii, represented by red, green, and blue curves.
In the upper panel of Fig. 1, we illustrate the evolution of the domain wall formed at time $t_{s}$ with a symmetry-breaking scale of $v=\sigma_{w}^{1/3}=10^{5}$ GeV. The formation time of the domain wall is determined by $t_{s}=M_{\rm Pl}/(3.32g_{*}^{1/2}T_{s}^{2})$, where $g_{*}(T)\simeq 106.75$ represents the effective degrees of freedom of the plasma, $M_{\rm Pl}=1.2\times 10^{19}$ GeV is the Planck mass, and the domain wall formation temperature is assumed to be $T_{s}=v$. The initial conditions $\bar{L}(t_{s})=0.01t_{s}$ and $\bar{v}_{w}\to 0$ are adopted to solve Eqs. (2.7). We further assume a vacuum pressure $V_{\rm bias}=(10\leavevmode\nobreak\ \rm GeV)^{4}$, with the onset of the bias potential occurring at $t_{\rm on}=0.01t_{\rm ann}$.
The red curves depict the evolution of $\bar{L}(t)$, while the blue curves represent $\bar{v}_{w}$. To emphasize the role of chopping, the parameter $c_{w}$ is set to $0.81$ and $0.61$ for the solid and dashed curves, respectively. We observe that the domain wall rapidly expands to a superhorizon size, $\bar{L}\sim 6t$, driven by curvature acceleration resulting from the wall tension. Subsequently, the correlation length of the wall decreases due to the Hubble friction, chopping effects, and friction from the bias potential. As the cosmic time approaches $t_{\rm ann}$, $\bar{L}(t)$ undergoes a significant reduction, re-entering the horizon before the domain wall annihilates at $t_{\rm ann}$. Initially, the domain walls are generated as a radiation gas, characterized by $\bar{v}_{w}=1$ and $\omega=1/3$ (see Eq. (2.6)). This behavior is consistent with the second equation in the VOS model (Eq. (2.7)), where setting $\bar{L}\to 0$ results in infinite curvature acceleration $k_{w}/\bar{L}\to\infty$. To ensure a finite initial velocity, $\bar{v}_{w}=1$ is imposed for the wall. However, due to resolution limitations, simulations do not explicitly capture this behavior [48]. Additionally, we note that at early times, chopping between walls can lead to an increase in $\bar{L}$, indicative of wall mergers. After reaching a maximum, chopping effectively reduces the wall size.
We observe that the domain wall rapidly expands to a superhorizon size, $\bar{L}\sim 6t$, driven by curvature acceleration resulting from the wall tension. Subsequently, the correlation length of the wall decreases due to Hubble friction, chopping effects, and friction from the bias potential. As the cosmic time approaches $t_{\rm ann}$, $\bar{L}(t)$ undergoes a significant reduction, re-entering the horizon before the domain wall annihilates at $t_{\rm ann}$. Initially, the domain walls are generated as a radiation gas, characterized by $\bar{v}_{w}=1$ and $\omega=1/3$ (see Eq. (2.6)). This behavior is consistent with the second equation in the VOS model (Eq. (2.7)), where setting $\bar{L}\to 0$ results in infinite curvature acceleration $k_{w}/\bar{L}\to\infty$. To ensure a finite initial velocity, $\bar{v}_{w}=1$ is imposed for the wall. However, due to resolution limitations, simulations do not explicitly capture this behavior [48]. Additionally, we note that at early times, chopping between walls can lead to an increase in $\bar{L}$, indicative of wall mergers. After reaching a maximum, chopping effectively reduces the wall size.
It is evident that the chopping effect should vary with time. However, the simulations in Ref. [48] assume a time-independent chopping parameter $c_{w}=0.81$, leading to the conclusion that the domain wall eventually enters the scaling regime. The chopping effect between walls can result in the efficient production of closed configurations, as confirmed by the simulations in Ref. [49]. The generation of closed domain walls could, in turn, suppress the chopping effects between walls. Once closed domain walls form at a superhorizon scale, the onset of the scaling regime for the walls could be hindered by the reduced chopping effect. In fact, the closed domain walls might never enter the scaling regime [43], contrary to previous expectations. To explore this, we adopt the following formula to describe the evolution of a spherical domain wall by neglecting the chopping effect [50]
$\ddot{\chi}+\left(4-3a^{2}\dot{\chi}^{2}\right)H\dot{\chi}+\frac{2}{a^{2}\chi}\left(1-a^{2}\dot{\chi}^{2}\right)=-\left(\frac{V_{\rm bias}}{\sigma_{w}}-6\pi G\sigma_{w}\right)\frac{\left(1-a^{2}\dot{\chi}^{2}\right)^{3/2}}{a},$ (2.10)
where the dot above a variable denotes a time derivative $d/dt$, and $\chi(t)$ is related to the curvature radius of the spherical wall by $\bar{L}(t)=a\chi(t)$, with the scale factor $a=(t/t_{i})^{1/2}$ (where $t_{i}$ denotes the initial time). Using $\rho_{w}=\mathcal{A}\sigma_{w}/t$ and $\rho_{\rm tot}=3H^{2}/(8\pi G)=3/(32\pi Gt^{2})$, the energy fraction of the domain wall is given by
$f_{w}(t)=\frac{\rho_{w}}{\rho_{\rm rad}}\simeq\frac{\rho_{w}}{\rho_{\rm tot}}=\frac{32\pi\mathcal{A}G\sigma_{w}t}{3}.$ (2.11)
Then, we obtain $G\sigma_{w}t=3f_{w}(t)/(32\pi\mathcal{A})$. Multiplying both sides of Eq. (2.10) by $t_{\rm ann}$, we derive the dimensionless form of Eq. (2.10) as
$\ddot{\tilde{\chi}}+\left(4-3a^{2}\dot{\tilde{\chi}}^{2}\right)\tilde{H}\dot{\tilde{\chi}}+\frac{2}{a^{2}\tilde{\chi}}\left(1-a^{2}\dot{\tilde{\chi}}^{2}\right)=-\left(C_{d}\mathcal{A}-\frac{9f_{w}(t_{\rm ann})}{16\mathcal{A}}\right)\frac{\left(1-a^{2}\dot{\tilde{\chi}}^{2}\right)^{3/2}}{a},$ (2.12)
[…]
In the lower panel of Fig. 1, we plot the trajectories of the curvature radius for different initial radii, represented by red, green, and blue lines. We observe that well before the horizon crossing of the domain wall, the wall’s radius scales as $\bar{L}(t)\propto a(t)$, indicating a superhorizon size above the horizon line $\bar{L}(t)=t$. As time progresses, the pressure from the bias potential becomes significant, causing the radius trajectory to cross the horizon line at a time $t_{e}$. If $t_{e}\lesssim t_{\rm ann}$, the wall network enters the horizon before significant annihilation occurs. For the case where $t_{e}\gtrsim t_{\rm ann}$, as denoted by the blue trajectory, the annihilation occurs at the horizon crossing time $t_{e}$, rather than at $t_{\rm ann}$, leading to the phenomenon of later annihilation [43]. Thus, we confirm that for closed domain walls with negligible chopping effects, the trajectory of the curvature radius remains at a superhorizon scale and does not enter the scaling regime ($\bar{L}(t)\simeq t$) until shortly before annihilation starts.